# HPC Alongside User-space Kubernetes

### v1

@@ III-C Hybrid Demonstration of Environment Capabilities
We ran 1000 training samples of LAMMPS using the Flux Framework workload manager with randomly chosen problem sizes ranging between 1 and 8 for each of parameters x, y, z to train three regression models served by an online machine learning server. We then ran 250 jobs of the same nature, but held out the actual wall time (Y) of the run and asked the server for a predicted value, saving the value for each request. While the specific results of the models are not a contribution of this paper, we can report that the paradigm to run HPC simulations with a workload manager alongside a Usernetes-owned service led to a successful outcome, with three different models for each of linear, bayesian regression, and passive aggressive regression. As we are primarily interested in the hybrid environment setup, we direct the interested reader to our results repository for diagrams of the models [73].
@@ IV Discussion
Despite the network performance not being entirely equal between Usernetes and “bare metal” setups, we were pleasantly surprised that the LAMMPS application had similar performance, and that differences in means were only a single order of magnitude, as opposed to 2x slower, which is what we saw in a similar virtual machine setup that used ethernet as the underlying network [64]. We believe this difference comes down to the likelihood that this particular build of LAMMPS is CPU bound at this smaller size, and using the elastic fiber adapter (EFA) vs. standard ethernet. The EFA verbs devices that are mounted into the container ignore the kernel entirely, and thus are not subject to the additional filtering of the TAP device. Messages over IBVerbs bypass that entirely, as it’s a direct message to the NIC. In practice, this means that messages over TCP might be subject to the extra processing done by the TAP device, but RDMA is not. The loss in performance we believe is due to other factors that warrant further work. A first area of investigation is if the Kubernetes headless Service that provides each of the Pods running Flux with a hostname. This hostname needs to be discovered by traversing a local lookup table, and this might be adding overhead. We also believe the other components of the software stack included with Flux or associated MPI libraries might not be bypassing slirp4netns. We anticipate further work in this area to continue to improve upon Usernetes’ ability to offer low latency.
@@
While the paradigm of an HPC workload manager running alongside infrastructure like Kuberneters is new, we can anticipate several exciting, possible future developments that will further support this setup. First, Kubernetes has support for custom schedulers. A current issue with the setup described here is that each of Usernetes and Flux Framework has their own scheduler, and we handled this issue by scheduling Usernetes within a Flux instance. The was worked around largely by requesting jobs that leave room for the other scheduler, however a more ideal approach would use a single scheduler that can account for scheduling to both resources, and this would be possible with a custom scheduler framework. Our group has done work in this space [92] and anticipates working on this. A second area for development that is specific to Flux Framework is to make it easy to submit workloads with logic that can orchestrate between the two environments. For example, if we take our hybrid machine learning prototype and extend it to be across node sizes and types (e.g., with or without GPU), we would want to use a Flux batch job with nested brokers that each control a scheduler for each group of resources. This is one of the elegant design features of Flux that was not used here, and we believe would immensely contribute to this design.
@@
We learned a great deal about good practices for deploying this kind of architecture in the AWS EC2 cloud, and specifically, attributes related to running user space Kubernetes alongside a workload manager on bare metal. First, while Pod resource requests and limits can be set to ensure a one to one mapping of Pods to nodes, and while the number of processes and cores might appear to be consistent between the virtual machine and a deployed Pod by using a tool like hwloc [93], this does not mean that the pod is able to fully utilize all resources. While we thought that a Pod in “Burstable” mode [94] could go over the resource limit, in practice this didn’t turn out to be the case. We discovered that setting a resource limit of CPU didn’t limit the CPU count, but did set a maximum for the percentage CPU cycles allowed, which can reduce an application’s CPU utilization from 99.9% down to 40-60%, which can have huge implications for performance. We encourage developers to only specify the EFA device in the resource requests and limits, and to use affinity to map one pod per node without impacting performance.
@@
Next, it’s important to be aware that systemd services can silently set limits on the workload manager’s ability to utilize memory or other system resources. If issues arise, we suggest investigating this setup. And finally, while outside of the scope of this work, we believe that other approaches that bypass the OS for an application to communicate with a network device should be explored, including (but not limited to) Infiniband and other network drivers.
@@
We also point to this project as a converged computing success in terms of collaboration. Developers from the HPC community, the authors of this paper, successfully collaborated with cloud native projects to push forward an important update to the Usernetes project that will likely further empower both sides. In the same thread, this work is an example of the power of small numbers of people. The Usernetes project is maintained, at most, by a handful of developers, and the authors of this paper were also small in numbers. However, through communication and advocacy, a novel paradigm was imagined and developed, and we believe it can have an impact on the larger HPC technology space. We encourage the larger community to also pursue ideas they believe in, and not be discouraged by their own sense of importance or numbers.

### v2

@@ III-C Hybrid Demonstration of Environment Capabilities
We ran 1000 training samples of LAMMPS using Flux with randomly chosen problem sizes ranging between 1 and 8 for each of parameters x, y, z to train three regression models served by an online machine learning server. We then ran 250 jobs of the same nature, but held out the actual wall time (Y) of the run and asked the server for a predicted value, saving the value for each request. While the specific results of the models are not a contribution of this paper, we can report that the paradigm to run HPC simulations with a workload manager alongside a Usernetes-owned service led to a successful outcome, with three different models for each of linear, bayesian regression, and passive aggressive regression. As we are primarily interested in the hybrid environment setup, we direct the interested reader to our results repository for diagrams of the models [51].
@@ IV Discussion
Despite the network performance not being equal between Usernetes and “bare metal” setups, we were surprised that the LAMMPS application had similar performance, and that differences in means were only a single order of magnitude, as opposed to 2×2$\times$2 × slower, which is what we saw in a similar virtual machine setup that used ethernet as the underlying network [44]. We believe this difference comes down to the likelihood that this particular build of LAMMPS is CPU bound at this smaller size, and using the elastic fabric adapter (EFA) vs. standard ethernet. The EFA verbs devices that are mounted into the container ignore the kernel entirely, and thus are not subject to the additional filtering of the TAP device. Messages over IBVerbs reach the NIC directly, implying that messages over TCP might be subject to the extra processing by the TAP device, but RDMA is not. The loss in performance may be due to other factors that warrant further study. A first area of investigation is the Kubernetes headless Service that provides each of the Pods running Flux with a hostname. This hostname needs to be discovered by traversing a local lookup table, and this might be adding overhead. It is also possible that other components of the software stack included with Flux or associated MPI libraries might not be bypassing slirp4netns. We anticipate further work in this area to continue to improve upon Usernetes’ ability to offer low latency.
@@
While the paradigm of an HPC workload manager running alongside infrastructure like Kuberneters is new, we can anticipate several exciting, possible future developments that will further support this setup. First, Kubernetes has support for custom schedulers. A current issue with the setup described here is that each of Usernetes and Flux Framework has their own scheduler, and we handled this issue by scheduling Usernetes within a Flux instance. The was worked around largely by requesting jobs that leave room for the other scheduler, however a more ideal approach would use a single scheduler that can account for scheduling to both resources, and this would be possible with a custom scheduler framework. Our group has done work in this space [73, 11] and will extend the work further. A second area for development that is specific to Flux Framework is to make it easy to submit workloads with logic that can orchestrate between the two environments. For example, if we take our hybrid machine learning prototype and extend it to be across node sizes and types (e.g., with or without GPU), we would want to use a Flux batch job with nested brokers that each control a scheduler for each group of resources. This is one of the core design features of Flux that was not used here, and we believe it would contribute to the flexibility and efficiency of the design.
@@
We learned a great deal about good practices for deploying this kind of architecture in the AWS EC2 cloud, and specifically, attributes related to running user space Kubernetes alongside a workload manager on bare metal. First, while Pod resource requests and limits can be set to ensure a one to one mapping of Pods to nodes, and while the number of processes and cores might appear to be consistent between the virtual machine and a deployed Pod by using a tool like hwloc [74], this does not mean that the pod is able to fully utilize all resources. While we thought that a Pod in “Burstable” mode [75] could go over the resource limit, in practice this didn’t turn out to be the case. We discovered that setting a resource limit of CPU did not limit the CPU count, but did set a maximum for the percentage CPU cycles allowed, which can reduce an application’s CPU utilization from 99.9% down to 40-60%, which can have huge implications for performance. We encourage developers to only specify the EFA device in the resource requests and limits, and to use affinity to map one pod per node without impacting performance.
@@
Next, it is important to be aware that systemd services can silently set limits on the workload manager’s ability to utilize memory or other system resources. If issues arise, we suggest investigating this setup. And finally, while outside of the scope of this work, we believe that other approaches that bypass the OS for an application to communicate with a network device should be explored, including (but not limited to) Infiniband and other network drivers.
@@
We also point to this project as a converged computing success in terms of collaboration. Developers from the HPC community, the authors of this paper, successfully collaborated with cloud native projects to push forward an important update to the Usernetes project that will likely further empower both sides. In the same vein, this work is an example of the power of small numbers of people. The Usernetes project is maintained, at most, by a handful of developers, and the authors of this paper were also small in numbers. However, through communication and advocacy, a novel paradigm was imagined and developed, and we believe it can have an impact on the larger HPC technology space. We encourage the larger community to also pursue ideas they believe in, and not be discouraged by their own sense of importance or numbers.
Vanessa Sochat is a Computer Scientist in Livermore Computing. Her work focuses on container technologies and high performance computing convergence in cloud. She maintains the Flux Operator and Fluence custom scheduler as part of the Flux Framework, and leads an ISCP for User-space Kubernetes. She collaborates closely with industry and several open source communities. She received her Ph.D and M.S. in Biomedical Informatics from Stanford University, and her B.A. from Duke University.
David Fox is a Computer Scientist in the Livermore Computing division at Lawrence Livermore National Laboratory, where his work involves, HPC, Cloud and Storage systems. He is a leader for the HPC Cluster Engineer Academy. He holds a B.S. in Electrical Engineering from Brigham Young University.
Daniel Milroy is a Computer Scientist at the Center for Applied Scientific Computing. His research focuses on graph-based scheduling and dynamic resource management for high-performance computing (HPC) and cloud-converged environments, and he collaborates closely with industry and academia. Dan is a developer of the Fluxion directed graph-based scheduler in the Flux Framework and leads a thrust of the Fractale LDRD. Dan holds a Ph.D. and M.S. in Computer Science from the University of Colorado Boulder, and an A.B. in Physics from the University of Chicago.